Compute the center of GL(3.F) Also find order of GL(3,3)
1
Expert's answer
2020-06-16T16:35:59-0400

"GL(3,F)=\\=\left\{A=\begin{pmatrix}
 a_{11} & a_{12}&a_{13} \\
 a_{21} & a_{22}&a_{23} \\
 a_{31} & a_{32}&a_{33} \\
\end{pmatrix}, 
a_{ij}\in F, detA=1\right\}"

The center is

"Z=\left\{z\in GL(3,F):\forall A\in GL(3,F):zA=Az\right\}\\
z=\begin{pmatrix}
 b_{11} & b_{12}&b_{13} \\
 b_{21} & b_{22}&b_{23} \\
 b_{31} & b_{32}&b_{33} \\
\end{pmatrix}\\
zA=\begin{pmatrix}
 b_{11} & b_{12}&b_{13} \\
 b_{21} & b_{22}&b_{23} \\
 b_{31} & b_{32}&b_{33} \\
\end{pmatrix}\cdot\begin{pmatrix}
 a_{11} & a_{12}&a_{13} \\
 a_{21} & a_{22}&a_{23} \\
 a_{31} & a_{32}&a_{33} \\
\end{pmatrix}\\
Az=\begin{pmatrix}
 a_{11} & a_{12}&a_{13} \\
 a_{21} & a_{22}&a_{23} \\
 a_{31} & a_{32}&a_{33} \\
\end{pmatrix}\cdot \begin{pmatrix}
 b_{11} & b_{12}&b_{13} \\
 b_{21} & b_{22}&b_{23} \\
 b_{31} & b_{32}&b_{33} \\
\end{pmatrix}"

We get the system

"a_{21}b_{12}+a_{31}b_{13}=a_{12}b_{21}+a_{13}b_{31}\\
a_{11}b_{21}+a_{21}b_{22}+a_{31}b_{23}=a_{21}b_{11}+a_{22}b_{21}+a_{23}b_{31}\\
a_{11}b_{31}+a_{21}b_{32}+a_{31}b_{33}=a_{31}b_{11}+a_{32}b_{21}+a_{33}b_{31}\\
a_{12}b_{11}+a_{22}b_{12}+a_{32}b_{13}=a_{11}b_{12}+a_{12}b_{22}+a_{13}b_{32}\\
a_{12}b_{21}+a_{32}b_{23}=a_{21}b_{12}+a_{23}b_{32}\\
a_{12}b_{31}+a_{22}b_{32}+a_{32}b_{33}=a_{31}b_{12}+a_{32}b_{22}+a_{33}b_{32}\\
a_{13}b_{11}+a_{23}b_{12}+a_{33}b_{13}=a_{11}b_{13}+a_{12}b_{23}+a_{13}b_{33}\\
a_{13}b_{21}+a_{23}b_{22}+a_{33}b_{23}=a_{21}b_{13}+a_{22}b_{23}+a_{23}b_{33}\\
a_{13}b_{31}+a_{23}b_{32}=a_{31}b_{13}+a_{32}b_{23}"

Since "detA\neq0" , the system has solutions if

"a_{11}\neq0, a_{22}\neq0, a_{33}\neq0, a_{ij}=0, i\neq j"

"z=\begin{pmatrix}
 a_{11} &0&0 \\
 0 &a_{22}&0\\
0&0&a_{33}
\end{pmatrix}"

Order "GL(3,Z_3)"

"(p^n-p^{n-1})(p^n-p^{n-2})(p^n-p^{n-3})...=\\
=(3^3-3^2)(3^3-3)(3^3-1)=18\cdot24\cdot26=11232"